OpenStudy (anonymous):

-6x-63= -75 What is the value of x

jigglypuff314 (jigglypuff314):

-6x-63= -75 add 63 to both sides +63 +63 -6x = -12 multiply both sides by (-6) /(-6) /(-6) x = (-12)/(-6) = 12/6 =

jigglypuff314 (jigglypuff314):

no... x = 12/6 = because negative divided by negative equals positive

OpenStudy (charlotte123):

x = 2 :D @jigglypuff314 12/6 = 2, so u were right ^_^ http://www.wolframalpha.com/input/?i=-6x-63%3D+-75+